a car slows down at -5.00m/s squared comes to a stop after 15 m what was the initial speed of the car.


1
Expert's answer
2022-02-22T10:30:28-0500

The distance traveled by a car to the rest

"d=\frac{v_i^2}{2a}"


Hence, the initial speed of a car

"v_i=\sqrt{2ad}=\sqrt{2*5.00*15}=12.2\;\rm m/s"